Hydrosulfuric acid, also known as hydrogen sulfide, is a diprotic acid. Its two-stage ionization is shown below:

H2S(aq) ⇆ H + + HS- (aq) Ka1 = 5.70x10-8

HS- (aq) ⇆ H + + S2- (aq) Ka2 = 1.0x 10-9

a. Calculate the concentration of HS- ion in a 0.222 M H2S solution.

b. Determine the pH of the solution.

c. Determine the S2- concentration.

Please write out all answers step-by-step :-)
Answer #1


Tofind thepH of hy drosulfuric aci d (HS) Given that H,S(aq) 근 H + + HS-(aq) Kal-5.70x10-8 HS-(aq) 근 H++52-(aq)Ka-10×10-9 0.2

Tofindthe S For the secondreaction HS-(aq) 근 H+4S-(aq)Ka=1.0x10-9 TH ES s*]=10×10-9 M Hence the HS 0.000112M pH = 3.95 [S2-]=